Calculate the power developed to lift a 8 kg mass to a height 2.4 meters in 6 seconds.



1
Expert's answer
2022-02-08T09:57:23-0500

"P=\frac At=\frac{Fh}t=\frac{mgh}t,"

"P=32~W."